OpenStudy (anonymous):

I am working on vectors but I don't know how to get the x and y components of a velocity vector. For example, a car is heading northwest at 9.00m/s. The components are (-6.36m/s)x and (6.36m/s)y for this particular vector. I understand that the x component is negative, but I don't understand how they got these numerical values.

OpenStudy (baru):

remember this x=V cos\(\theta\) y=V sin\(\theta\) where \(\theta\) is the angle made with the positive x axis

OpenStudy (baru):

V is the magnitude of the vector

OpenStudy (baru):



OpenStudy (unklerhaukus):